OpenStudy (anonymous):

What line is perpendicular to y=-3x+1; and also passes through (2,2)?

OpenStudy (anonymous):

clearly the given line is of the form y=mx +c hence the slope of the given line =-3 hence slope of the line perpendicular to it is 1/3 hence the eq of the reqd line passing through (2,2) is given by (y-2)=1/3 *(x-2)

OpenStudy (anonymous):

3y-6=x-2 x-3y+4=0

OpenStudy (anonymous):

@matricked is it possible to put it into y=mx+b?

OpenStudy (anonymous):

here b or c irrelevant

OpenStudy (anonymous):

for the line given by y=mx+c m is the slope and c is the y-intercept

OpenStudy (anonymous):

ok, thanks very much! @matricked

OpenStudy (anonymous):

mention not dear